Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 424     Entry time: 30 Oct 2023, 07:10
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 7.8keV, ~2.5mj, no beam before 1:15am
  • One undulator was not moving, expert was called

Optical laser delivery

  • ns laser for particle imaging

Achievements/Observations

  • RGA used for quantitative gas BG analysis
  • Collected more data from 55 nm Ag cube
  • Collected some sucrose data
  • Collected gold bipyramids data

Issues

  • Could not get sucrose droplets delivered to the interation point while injecting He at the bottom of the skimmer. As soon as it was enough He to decrease N, droplets desappeared.
  • Still some questionmarks regarding the gas flows through the skimmer section into the chamber. More tests in the lab necessary.
ELOG V3.1.4-7c3fd00